# Characterizing the Role of Similarity in the Property Inferences of Language Models

### v1

@@ Abstract
Characterizing the Role of Similarity in the Property Inferences of Language Models
@@ 1 Introduction
Categories are fundamental to human semantic cognition. Our knowledge of categories allows us to draw everyday inferences, and one such category-based inference is property inheritance, where properties are projected from a category to its members. For example, if we learn that dogs have the T9 hormone, we can reasonably assume that corgis also have the T9 hormone. Psychologically, an obvious and previously popular explanation for property inheritance is that it is the natural consequence of our minds organizing categories hierarchically into taxonomies (Collins and Quillian, 1969; Glass and Holyoak, 1974; Murphy, 2002).
At the same time, empirical evidence has called this assumption into question. Specifically, Sloman (1998) showed that humans were highly sensitive to the similarity of the categories when performing property inheritance. That is, they were more likely to project novel properties from birds to robins (a more typical bird) than from birds to penguins (a less typical bird), despite them agreeing that both were members of the bird category. Sloman concluded that instead of explicitly using stored taxonomies, humans might simply be computing categorical similarities to demonstrate inheritance-compatible behavior; if so, a property is more likely to be shared between categories insofar as the similarity between them is high enough.
@@
Debates about the mechanisms that underlie human property inheritance offer an exciting avenue to analyze conceptual organization and use in language models (LMs). First, due to the close connection of a word to its conceptual meaning (Murphy, 2002; Lupyan and Lewis, 2019; Lake and Murphy, 2021), property inheritance provides an opportunity to diagnose meaning-sensitivities in LMs (Piantadosi and Hill, 2022; Bender and Koller, 2020). Next, similarity effects like the ones found by Sloman (1998) are related to the notion of content effects shown by humans and LMs on logical reasoning tasks (Wason, 1968; Evans, 1989; Lampinen et al., 2024). That is, the presence of similarity effects goes against the idea that humans use abstract taxonomic principles to perform property inheritance (where every category member is equally likely to inherit the property). Finally, the principles of taxonomy vs. similarity offer two clear hypotheses about what mechanisms could govern property inheritance behavior in a system, allowing us to test them in LMs. This allows us to go beyond prior work investigating property inheritance in LMs (Misra et al., 2023), which has largely focused on their behavior, leaving open questions about the underlying mechanisms.
In this work, we empirically investigate the roles of taxonomic relations and categorical similarities on LMs’ property inheritance behavior, using both behavioral and causal interpretability methods on four LMs. First, we behaviorally analyze LMs’ sensitivity to taxonomic relations when performing property inheritance. We demonstrate that noun similarity, known to correlate with human property inheritance behavior (Sloman, 1998), also correlates strongly with property inheritance judgments in language models, and explains many of the false positives and false negatives. Then, we causally localize property inheritance to specific subspaces of the activation space via distributed alignment search (DAS; Geiger et al., 2024b). DAS is a causal interpretability method that aligns subspaces in the network with a hypothesized causal model of the task at hand. By training and evaluating DAS on controlled subsets of the data, we show that the subspaces responsible for property inheritance reflect both taxonomic and similarity features. Together, these results indicate that LMs are susceptible to non-trivial content effects rooted in noun similarities, and moreover that taxonomy and similarity are fundamentally entangled in model representations.
@@ 2 Conceptual Organization in LMs
In this vein, a number of recent works have investigated LMs’ property inheritance behavior—either by purely behavioral analyses (Misra et al., 2023; Wu et al., 2023a), or by manipulating LMs’ internal representations to edit their taxonomic knowledge (Powell et al., 2024; Cohen et al., 2024), and measuring the implications of the edit. Importantly, these works assume that changes to taxonomic membership (by editing or introducing novel concepts) must be directly reflected in the LMs’ behavior, without necessarily considering the impact of categorical similarity. Grounding our investigation in empirical observations about humans (Sloman, 1998), we abstract away from this assumption, and explicitly consider the interplay between similarity and taxonomic relations by characterizing it using both behavioral and causal analysis methods.
@@ 3.1 Data
Since we are explicitly analyzing the effect of similarity in LMs’ property inheritance behavior, we include similarity measures as a central component for deriving pairs of related categories that are not taxonomically related. We perform this sampling as follows: for each superordinate category $C$, we construct the negative sample space $\mathcal{N}$ as the set of items that are outside that category—e.g., for bird, this could mean non-birds such as zebra, sofa, etc. We then compute the similarity of each item in $\mathcal{N}$ to the superordinate category using a given similarity measure. Finally, if $k$ is the number of category members of $C$, then we sample the top $\nicefrac{{k}}{{2}}$ and bottom $\nicefrac{{k}}{{2}}$ concepts from $\mathcal{N}$, based on the similarity values. This way, for each superordinate category, we have $k$ category members, and $k$ non-category members, with an even split between high and low-similarity non-taxonomic items.
@@ 3.2 Similarity scores
Table 2: Behavioral sensitivities of LMs across both similarity types (Word-Sense and SPoSE). TS: Taxonomic Sensitivity; PS: Property Sensitivity; MS: Mismatch Sensitivity; $\rho$: Spearman correlation with similarity.
@@ 4 Causal Investigation of Taxonomy and Similarity
Given that LMs perform property inheritance with high taxonomic sensitivity, we now use causal interpretability tools (Mueller et al., 2024a; Geiger et al., 2024a) to localize this behavior. We then investigate (i) whether the subspaces that are causally responsible for inheritance perform inductive generalization in a manner that is better explained by the similarity of the noun concepts, rather than their taxonomic relationship; and (ii) whether taxonomic information and noun similarity are fundamentally entangled in the model representations. We investigate both questions using distributed alignment search (DAS; Geiger et al., 2024b), a causal interpretability method which been previously used to localize and mechanistically explain syntax-sensitive behaviors (Arora et al., 2024) and arithmetic abilities (Wu et al., 2023b) in LMs. To avoid needing to search over all possible subspaces of an activation vector manually, we employ the Boundless DAS variant (Wu et al., 2023b) implemented in pyvene (Wu et al., 2024), to learn the appropriate subspace.
@@
We use DAS rather than other causal interpretability techniques [^7] For example, those based on sparse autoencoders (Bricken et al., 2023; Marks et al., 2024; Huben et al., 2024) or neurons (Vig et al., 2020; Finlayson et al., 2021). because it allows us to investigate to what extent a specific hypothesized causal model is being implemented by the neural network. Furthermore, because DAS isolates a specific subspace corresponding to a given hypothesized causal variable, the subspace can be evaluated in novel contexts post-hoc to better characterize its functional role. We leverage this to evaluate whether subspaces that are sensitive to taxonomic relationships are also sensitive to similarity-based but non-taxonomic relationships. This type of experimental design involving ambiguous training sets with separate test and generalization sets has been used to characterize the inductive biases of LMs (McCoy et al., 2019; Si et al., 2023; Mueller et al., 2024b), though this has typically been evaluated behaviorally rather than at the level of internal mechanisms. There is a limited but growing literature validating how well discovered mechanisms (or explanations thereof) generalize to novel types of contexts (Geiger et al., 2020; Huang et al., 2023, 2024, i.a.). Geiger et al. (2020) in particular used this technique to causally analyze lexical entailment in LMs, whereas we target the role of taxonomy vs. similarity.
@@
Given an LM, a hypothesized causal graph, and counterfactual input pairs that target a specific variable in the causal graph, DAS learns a rotation for the activations of a given submodule, and then performs counterfactual interventions in this rotated space. The intuition is that the features of interest may not necessarily be aligned to the bases of the original activation space (Hinton et al., 1986; Smolensky, 1986; Elhage et al., 2022; Mueller et al., 2024a), rendering neuron-based interventions insufficient for capturing the target variable. We refer readers to Geiger et al. (2024b) for details.
After learning the rotation, we compute the interchange intervention accuracy (IIA) on a held-out test set. This scores how frequently the learned subspace performs as expected given the hypothesized causal graph. In our experiments, we hypothesize that a given subspace will be sensitive to taxonomic relationships—if this is true, then passing an activation into this subspace from taxonomically unrelated inputs should yield a variable output of False. However, if we intervene on this subspace passing in activations from an alternate input where the nouns are taxonomically related (while holding all other activations in the model constant), then the variable’s output should flip to True. We illustrate the intuition and the causal graph that we hypothesize the LMs implement in Figure 3. By only changing the premise and conclusion nouns across inputs (e.g., birds, robins, orchids), and keeping the rest of input constant, we effectively target the node “B is a kind of A” in Figure 3. We target this node since our focus is on taxonomic relations rather than property binding—where the goal is to check if LMs have appropriately bound the property with the noun concept—which has been investigated in other work (Feng and Steinhardt, 2024).
We train DAS on a subset of 3,000 stimuli and evaluate on the rest. Implementation details on DAS training are given in Appendix D. Next we describe variations of the training and test sets used in our experiments.
@@ 4.1 Results and Analysis
IIA values for the control experiment are always lower than for the balanced setting; thus, despite its expressivity, DAS can yield task-dependent results. High values for IIA occur in roughly the same positions for Ambiguous-Test and Ambiguous-Gen. There is a substantial drop from Ambigious-Test to Ambigious-Gen for all models and for every intervention position, showing that, when given ambiguous data, the learned rotations tend to rely on similarity rather than taxonomic relations. This was more substantial in some models (Mistral-7B) than in others (Gemma-2-9B). This effect is stronger for SPoSE negative examples than for Word-Sense negative examples, suggesting that LMs’ property inheritance behavior corresponds more closely with perceptual and conceptual properties (i.e., which SPoSE claims to have captured, see Kaniuth et al., 2024) than general distributional similarity (which aggregates over all possible contexts in which the noun concepts occur in text). Similar observations hold for the Unambigious experiment, with a drop in scores showing that, even when trained on data that is unambigiously consistent with taxonomic relations rather than similarity, the causal model is sensitive to similarity effects. In other words, similarity and taxonomic features appear to be fundamentally entangled.
@@
Table 3: Maximum IIA values for the Unambiguous experiment. Interventions are trained under the Balanced train set, and evaluated on both the Balanced test set (Bal) as well as the Ambiguous-Gen set (Gen). The drop on the generalization test set shows that the learned interventions rely non-trivially on similarity.
5 Discussion and Conclusions
Our goal in this paper was to characterize category-based inferences in LMs. We derived inspiration from experiments on humans which sought to understand how they transfer properties from a higher level category like mammals to lower level ones such dogs and cats (Collins and Quillian, 1969; Glass and Holyoak, 1974; Sloman, 1998)—i.e., show property inheritance. Findings from these experiments suggest that instead of relying purely on abstract category-membership principles (where all members of a category are equally likely to inherit properties), humans show graded behavior proportional to the similarity between the premise and the conclusion Sloman (1998).
Through behavioral and representational analyses methods across four LMs, we found that language models’ property inheritance behavior is sensitive to both taxonomic relations and similarity. Behaviorally, LMs were more likely to extend the property from the premise to the conclusion when they were taxonomically related, and at the same time, showed positive correlations with the similarity between them. This finding was further reinforced in our representational investigations, where we found taxonomy and similarity to be largely entangled in the LM subspaces causally responsible for property inheritance. Combined with the human experiments, our results cast doubt on previous works’ assumption of pure taxonomic principles in the context of property inheritance in LMs, without considering similarity (Misra et al., 2023; Powell et al., 2024; Wang et al., 2024).
Our findings also suggest interesting hypotheses for future human analyses. Importantly, Sloman (1998)’s experiments never tested humans on cases where the categories in a property inheritance setting did not share taxonomic relations (e.g., birds and bats). This makes it unclear if the principles of taxonomy and similarity are also not mutually exclusive for humans (Murphy et al., 2012)—e.g., it could be possible that both mechanisms are at play, where taxonomic relations demarcate the inference (i.e., whether the property will be inherited) and similarity modulates its strength. Future experiments could test this hypothesis and shed further light on the extent to which LMs and humans show similar sensitivities and biases during reasoning.
@@ 5 Discussion and Conclusions
6 Acknowledgments
A.M. is supported by a postdoctoral fellowship under the Zuckerman STEM Leadership program. K.M. acknowledges postdoctoral funding supported by NSF Grant 2139005 awarded to Kyle Mahowald. We also thank Katrin Erk for helpful discussion and feedback.

### v2

@@ Abstract
Characterizing the Role of Similarity in the
Property Inferences of Language Models
@@ 1 Introduction
Categories are fundamental to human semantic cognition. Our knowledge of categories allows us to draw everyday inferences; a prominent example of which is property inheritance, where properties are projected from a category to its members. For example, if we learn that dogs have the T9 hormone, we can reasonably assume that corgis also have the T9 hormone. In cognitive psychology, an obvious and previously popular explanation for property inheritance is that it is the natural consequence of our minds organizing categories hierarchically into taxonomies (Collins and Quillian, 1969; Glass and Holyoak, 1974; Murphy, 2002).
At the same time, empirical evidence has called this assumption into question. Sloman (1998) showed that humans were highly sensitive to the similarity of the categories when performing property inheritance; for example, they were more likely to project novel properties from birds to robins (a more typical bird) than from birds to penguins (a less typical bird), despite them agreeing that both were members of the bird category. Sloman concluded that instead of explicitly using stored taxonomies, humans might simply be computing categorical similarities to demonstrate inheritance-compatible behavior. In such a case, a property is more likely to be shared between categories insofar as the similarity between them is high enough.
@@
Debates about the mechanisms that underlie human property inheritance offer an exciting avenue to analyze conceptual organization and use in language models (LMs). First, due to the close connection of a word to its conceptual meaning (Murphy, 2002; Lupyan and Lewis, 2019; Lake and Murphy, 2021), property inheritance provides an opportunity to diagnose meaning-sensitivities in LMs (Piantadosi and Hill, 2022; Bender and Koller, 2020). Next, similarity effects like those found by Sloman (1998) are related to the notion of content effects shown by humans and LMs on logical reasoning tasks (Wason, 1968; Evans, 1989; Lampinen et al., 2024); the presence of similarity effects goes against the idea that humans rely on abstract taxonomic principles (where every category member would be equally likely to inherit the property). Finally, the principles of taxonomy vs. similarity offer two clear hypotheses about what mechanisms could govern property inheritance behavior in a system. This allows us to go beyond prior work investigating property inheritance in LMs (Misra et al., 2023), which has largely focused on their behavior, leaving open questions about the underlying mechanisms.
In this study, we empirically investigate the roles of taxonomic relations and categorical similarities on LMs’ property inheritance behavior, using both behavioral and causal interpretability methods on four LMs. First, we behaviorally analyze LMs’ sensitivity to taxonomic relations when performing property inheritance (§3.4). We demonstrate that noun similarity, known to correlate with human property inheritance behavior (Sloman, 1998), also correlates strongly with property inheritance judgments in language models, and explains many of the false positives and false negatives. Then, we causally localize property inheritance to specific activation subspaces (§4) using distributed alignment search (DAS; Geiger et al., 2024b). By training and evaluating DAS on controlled subsets of the data, we show that the subspaces responsible for property inheritance reflect both taxonomic and similarity features. Together, these results indicate that LMs are susceptible to non-trivial content effects rooted in noun similarities, and moreover that taxonomy and similarity are fundamentally entangled in model representations.
@@ 2 Conceptual Organization in LMs
To what extent is the similarity vs. taxonomy debate relevant to LMs? By definition, LMs learn representations from distributional statistics of tokens in context, so any reasoning they demonstrate is a result of distributional similarity. While technically true, this viewpoint may be an oversimplification: two words can be distributionally similar if they participate in antonymy (hot vs. cold), metonymy (car, wheel), hypernymy (robin, bird), or even if they share thematic relations (dog, bone). For property inheritance, however, it is only the hypernymy relation that plays a critical role (Murphy, 2002). Is it possibile that something structural like a taxonomy might arise through distributional statistics? It has been postulated that symbolic behavior can emerge as a result of sub-symbolic processes internal to a neural network (Smolensky, 1986, 1988; McClelland et al., 2010). Empirically, this has been shown in a number of recent works, e.g., Nanda et al. (2023) on LMs and modular arithmetic, and Feng et al. (2023) on LMs learning to simulate dynamic programming—both of which require operating over symbolic representations.
The above discussion raises multiple interrelated questions: when performing property inheritance, do LMs show sensitivity to taxonomic relations? If so, is their behavior graded, where properties are less likely to be projected to concepts that are taxonomically related to the higher-level category but low in similarity? Or, is it more binary, where there is no significant difference between how likely a property is projected from the higher-level category to its high- and low-similarity members? If their behavior is graded, then does it differ between non-category members that might share similarities with the higher level category (e.g., bird and giraffe) vs. those that do not (e.g., bird and chair)?
Related work
@@
In this vein, recent works have investigated LMs’ property inheritance behavior—either by purely behavioral analyses (Misra et al., 2023; Wu et al., 2023a), or by manipulating LMs’ internal representations to edit their taxonomic knowledge (Powell et al., 2024; Cohen et al., 2024), and measuring the implications of the edit. Importantly, these works assume that changes to taxonomic membership must be directly reflected in the LMs’ behavior, without necessarily considering the impact of categorical similarity. Grounding our investigation in empirical observations about humans (Sloman, 1998), we abstract away from this assumption, and explicitly consider the interplay between similarity and taxonomic relations by characterizing it using both behavioral and causal analysis methods.
@@ 3.1 Data
Because we are explicitly analyzing the effect of similarity in LMs’ property inheritance behavior, we include similarity measures as a central component for deriving pairs of related categories that are not taxonomically related. We perform this sampling as follows: for each superordinate category $C$, we construct the negative sample space $\mathcal{N}$ as the set of items that are outside that category—e.g., for bird, this could mean non-birds such as zebra, sofa, etc. We then compute the similarity of each item in $\mathcal{N}$ to the superordinate category using a given similarity measure. Finally, if $k$ is the number of category members of $C$, then we sample the top $\nicefrac{{k}}{{2}}$ and bottom $\nicefrac{{k}}{{2}}$ concepts from $\mathcal{N}$, based on the similarity values. This way, for each superordinate category, we have $k$ category members, and $k$ non-category members, with an even split between high and low-similarity non-taxonomic items.
@@ 3.2 Similarity scores
[…]
@@ 4 Causal Investigation of Taxonomy and Similarity
Given that LMs perform property inheritance with high taxonomic sensitivity, we now use causal interpretability tools (Mueller et al., 2024a; Geiger et al., 2024a) to localize this behavior. We then investigate (i) whether the subspaces that are causally responsible for inheritance perform inductive generalization in a manner that is better explained by the similarity of the noun concepts, rather than their taxonomic relationship; and (ii) whether taxonomic information and noun similarity are fundamentally entangled in the model representations. We investigate both questions using distributed alignment search (DAS; Geiger et al., 2024b), a causal interpretability method which has been used to localize and mechanistically explain syntax-sensitive behaviors (Arora et al., 2024) and arithmetic abilities (Wu et al., 2023b) in LMs. To avoid needing to search over all possible subspaces of an activation vector manually, we employ the Boundless DAS variant (Wu et al., 2023b) implemented in pyvene (Wu et al., 2024) to learn the appropriate subspace.
@@
We use DAS rather than other causal interpretability techniques [^7] For example, those based on sparse autoencoders (Bricken et al., 2023; Marks et al., 2024; Huben et al., 2024) or neurons (Vig et al., 2020; Finlayson et al., 2021). because it allows us to investigate to what extent a specific hypothesized causal model is being implemented by the neural network. Because DAS isolates a specific subspace corresponding to a given hypothesized causal variable, the subspace can be evaluated in novel contexts post hoc to better characterize its functional role. We leverage this to evaluate whether subspaces that are sensitive to taxonomic relationships are also sensitive to similarity-based but non-taxonomic relationships. This type of experimental design involving ambiguous training sets with separate test and generalization sets has been used to characterize the inductive biases of LMs (McCoy et al., 2019; Si et al., 2023; Mueller et al., 2024b), though this has typically been evaluated behaviorally rather than at the level of internal mechanisms. There is a limited but growing literature validating how well discovered mechanisms (or explanations thereof) generalize to novel types of contexts (Geiger et al., 2020; Huang et al., 2023, 2024, i.a.).
Given an LM $M$, a hypothesized causal graph $L$, and counterfactual input (source, base) pairs $(s,b)$ that target a specific intermediate variable $V$ in the causal graph, DAS learns a rotation $R$ for the activations of a given submodule, and then performs counterfactual interventions in this rotated space. This is done as follows: $L_{V}(s)$, the value of variable $V$ in $L$ when applied to source $s$, is patched into the corresponding variable of the base $b$,
$L_{V}(b)\leftarrow L_{V}(s),$
and then the final output of $L$ is computed. Translating this to operations on the low-level computation graph, a rotation $R$ is learned for activations in the network $M$ from a position $i$ to position $j$. We patch $R(M(s)_{i})$ into $M(b)_{j}$:
$M(b)_{j}\leftarrow R\bigl{(}M(s)_{i}\bigr{)},$
and then compute the output of $M$.
After learning the rotation, we compute the interchange intervention accuracy (IIA) on a held-out test set. This scores how frequently the learned subspace performs as expected given the hypothesized causal graph. Intuitively, IIA is defined as the proportion of $(s,b)$ pairs where patching in the rotated source activation into the base model $M$ causes $M$’s output to match the causal model $L$’s output after the intervention. It thus measures the degree of alignment between the neural network and the causal graph. The intuition for learning a rotation is that the features of interest may not necessarily be aligned to the bases of the original activation space (Hinton et al., 1986; Smolensky, 1986; Elhage et al., 2022; Mueller et al., 2024a), rendering neuron-based approaches insufficient for capturing the target variable. We refer readers to Geiger et al. (2024b) for details.
@@
In our experiments, we hypothesize that a given subspace will be sensitive to taxonomic relationships—if this is true, then passing an activation into this subspace from taxonomically unrelated inputs should yield a variable output of False. However, if we intervene on this subspace passing in activations from an alternate input where the nouns are taxonomically related (while holding all other activations in the model constant), then the variable’s output should flip to True. We illustrate the intuition and the causal graph that we hypothesize LMs implement in Figure 3. By only changing the premise and conclusion nouns across inputs (e.g., birds, robins, orchids), and keeping the rest of input constant, we effectively target the node “B is a kind of A” in Figure 3.888We target this node since our focus is on taxonomic relations rather than property binding—where the goal is to check if LMs have appropriately bound the property with the noun concept—which has been investigated in other work (Feng and Steinhardt, 2024).
We train DAS on a subset of 3,000 stimuli and evaluate on the rest. Implementation and training details are given in Appendix D. Next, we describe variations of the training and test sets used in our experiments.
@@ 4.1 Results and Analysis
IIA values for the control experiment are always lower than for the balanced setting; thus, despite its expressivity, DAS can yield task-dependent results (i.e., comparatively lower IIA values in the control setting establishes that DAS is telling us something about the LM’s property inheritance abilities). High values for IIA occur in roughly the same positions for Ambiguous-Test and Ambiguous-Gen. There is a substantial drop from Ambigious-Test to Ambigious-Gen for all models and for every intervention position, showing that, when given ambiguous data, the learned rotations tend to rely on similarity rather than taxonomic relations. Similar observations hold for the Unambigious experiment, with a drop in scores showing that, even when trained on data that is unambigiously consistent with taxonomic relations rather than similarity, the causal model is sensitive to similarity effects. In other words, similarity and taxonomic features appear to be fundamentally entangled.
5 Discussion and Conclusions
Our goal in this paper was to characterize category-based inferences in LMs. We derived inspiration from psychological experiments that have sought to understand how humans transfer properties from a higher level category like mammals to lower level ones such dogs and cats (Collins and Quillian, 1969; Glass and Holyoak, 1974; Murphy, 2002)—i.e., show property inheritance. Findings from these experiments show that instead of relying purely on abstract category-membership principles (where all members of a category are equally likely to inherit properties), humans show graded behavior proportional to the similarity between the categories Sloman (1998).
@@ 5 Discussion and Conclusions
Table 3: Maximum IIA values for the Unambiguous experiment. Interventions are trained under the Balanced train set, and evaluated on both the Balanced test set (Bal) as well as the Ambiguous-Gen set (Gen). The drop on the generalization test set suggests that the learned interventions rely non-trivially on similarity.
Through behavioral and representational analyses methods across four LMs, we found that language models’ property inheritance behavior is sensitive to both taxonomic relations and similarity. Behaviorally, LMs were more likely to extend the property from the premise to the conclusion when they were taxonomically related, and at the same time, showed positive correlations with the similarity between them. This finding was further reinforced in our representational investigations, where we found taxonomy and similarity to be largely entangled in the LM subspaces causally responsible for property inheritance. Our results cast doubt on previous works’ assumption of pure taxonomic principles with respect to property inheritance in LMs, without considering similarity (Misra et al., 2023; Powell et al., 2024; Wang et al., 2024). In the broader context of LMs’ reasoning behaviors, our results showcase yet another instance where LMs were susceptible to human-like content effects (Lampinen et al., 2024).
Our findings also suggest interesting hypotheses for future human analyses. Importantly, Sloman (1998)’s experiments never tested humans on cases where the categories in a property inheritance setting did not share taxonomic relations (e.g., birds and bats). This makes it unclear if the principles of taxonomy and similarity are also not mutually exclusive for humans (Murphy et al., 2012)—e.g., it could be possible that both mechanisms are at play, where taxonomic relations demarcate the inference (i.e., whether the property will be inherited) and similarity modulates its strength. Future experiments could test this hypothesis and shed further light on the extent to which LMs and humans show similar sensitivities and biases during reasoning. Overall, our work shows how tools from modern interpretability research, primarily used to isolate sub-structures in neural networks, may also be used to perform hypothesis testing of the abstractions encoded within them.
@@
Acknowledgments
A.M. is supported by a postdoctoral fellowship under the Zuckerman STEM Leadership program. K.M. acknowledges postdoctoral funding supported by NSF Grant 2139005 awarded to Kyle Mahowald. We also thank Martin Hebart for help on the THINGS dataset and the SPoSE embeddings, Atticus Geiger and Zhengxuan Wu for assistance with pyvene, and Katrin Erk and the anonymous reviewers for helpful discussion and feedback.